# Evaluating Large Language Models as Virtual Annotators for Time-series Physical Sensing Data

### v1

@@ Abstract.
Traditional human-in-the-loop-based annotation for time-series data like inertial data often requires access to alternate modalities like video or audio from the environment. These alternate sources provide the necessary information to the human annotator, as the raw numeric data is often too obfuscated even for an expert. However, this traditional approach has many concerns surrounding overall cost, efficiency, storage of additional modalities, time, scalability, and privacy. Interestingly, recent large language models (LLMs) are also trained with vast amounts of publicly available alphanumeric data, which allows them to comprehend and perform well on tasks beyond natural language processing. Naturally, this opens up a potential avenue to explore LLMs as virtual annotators where the LLMs will be directly provided the raw sensor data for annotation instead of relying on any alternate modality. Naturally, this could mitigate the problems of the traditional human-in-the-loop approach. Motivated by this observation, we perform a detailed study in this paper to assess whether the state-of-the-art (SOTA) LLMs can be used as virtual annotators for labeling time-series physical sensing data. To perform this in a principled manner, we segregate the study into two major phases. In the first phase, we investigate the challenges an LLM like GPT-4 faces in comprehending raw sensor data. Considering the observations from phase 1, in the next phase, we investigate the possibility of encoding the raw sensor data using SOTA SSL approaches and utilizing the projected time-series data to get annotations from the LLM. Detailed evaluation with four benchmark HAR datasets shows that SSL-based encoding and metric-based guidance allow the LLM to make more reasonable decisions and provide accurate annotations without requiring computationally expensive fine-tuning or sophisticated prompt engineering.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
@@ 1. Introduction
Interestingly, most recent large language models (LLMs) have also been trained with vast knowledge from publicly available data, potentially containing information from numeric data beyond the usual text data (Liu et al., 2023; Yao et al., 2022). This, in turn, opens up an avenue to investigate whether these LLMs can directly annotate the raw sensor data without using alternate modalities. To assess this, we first set up a detailed pilot study (Section 4) whereby we analyze whether the state-of-the-art (SOTA) LLMs like GPT-4 can quickly identify the labels (the activity classes) in a binary classification setup. This study shows that although GPT-4 can understand the accelerometer signatures, it fails to separate them across different classes even with a fixed number of bootstrap samples as example observations. Notably, some recent studies like (Liu et al., 2023) have already attempted to investigate how LLMs can be used with sensor data streams and reached similar conclusions.
@@
    Extending LLM’s Capability as Virtual Annotators of the Physical Sensing Data: Based on the findings from the pilot study, we develop an approach to encode the locomotive data in an embedding space that inherently learns the signatures of different classes (in a binary setup) and thus provides rich contextual information to the LLMs for annotating the accelerometer signatures with the latent embedding of the raw locomotive data. We explore self-supervised contrastive learning with two different setups for this purpose – (a) with time-domain signatures only and (b) with both time and frequency-domain signatures. To the best of our knowledge, this is the first work that helps LLMs enhance their capability for annotating locomotive data from the raw accelerometer with the corresponding human mobility classes without requiring any retraining.
@@
    Thorough Evaluation, Analysis, and Discussion: We perform a detailed investigation using four benchmark HAR datasets on how the encoded accelerometer data offers better reasoning for the LLM to provide accurate annotations with the increasing number of examples. Notably, with paid LLMs like ChatGPT-4, a mere performance assessment with accuracy as the key highlight is not sufficient as there are factors like the number of tokens and rate limits are restrictions that impact the final usage of such a service for annotating a dataset. Understanding this, we in this paper, also assess the time and cost of analysis, summarise the trade-offs, and discuss the advantages of using LLMs as a virtual annotator for time-series physical sensing data.
@@ 2.1. Using LLMs for Understanding Time-Series Data
Interestingly, works like (Liu et al., 2023) used the raw sensor data as an example; however, the idea did not show consistent performance with increasing classes. More recent works, like (Brown et al., 2020; Kim et al., 2024; Xue et al., 2022; Xu et al., 2023a) have attempted to provide more guidance and examples through in-context learning which allows the inclusion of more information within the prompt and does not necessarily need any fine-tuning of the LLM parameters. This approach also has the immense benefit of allowing the transfer of human expertise to the LLM (Xu et al., 2023a; Xie and Min, 2022). Furthermore, works like (Kim et al., 2024; Xu et al., 2023a) also show the importance of providing signal-processing (Xu et al., 2023a) or interpretation-based guidance (Kim et al., 2024) which allows the model to make predictions regarding the query sample. However, one key aspect that we observe in most of these aforementioned works is that in most cases, the input time series is from specialized sensors like medical (Kim et al., 2024) or temperature sensors (Xu et al., 2023a) for which expert guidance is possible. For example, the prompt can provide what a typical high heart rate means and how it can be observed in the ECG signal. However, with more generalized sensors like accelerometers, it is hardly possible to provide some expert guidance directly, given the inherent nature of the sensor data itself. To mitigate this, works like (Jin et al., 2024) have discussed the idea of using time-series encoders, albeit with a proposal to align the time-series features with language-specific representations, which in turn would need further fine-tuning of the LLM. In this study, however, we explore using a time-series encoder to provide more context to the LLM without fine-tuning. Based on this objective, we assess whether an LLM like GPT-4 can be used as a service to obtain annotations for the unlabeled accelerometer data for HAR.
@@ 3. Datasets Used for the Analysis
UCI HAR (Reyes-Ortiz Jorge, 2012) 164718 47033 Waist Standing & Walk upstairs 50 Hz
HHAR (Blunck Henrik and Anind, 2015) 3020605 521516 Arm Standing & Stairs up 50$\sim$200 Hz
@@
    UCI HAR – The UCI HAR dataset (Reyes-Ortiz Jorge, 2012) contains inertial data for 6 different activities (Laying, Sitting, Standing, Walking, Walking downstairs and Walking upstairs). These activities were performed by wearing a smartphone around their waist. The IMUs have a sampling rate of 50Hz. For our setup, we have chosen “Standing” and “Walking upstairs” as the activity classes and subjects 1, 3, 5, 7, 9, 11, 13, 15 and 17 as the test subjects whereas, all other remaining subjects across all activity classes are part of the training set.
@@
    HHAR – The HHAR dataset (Blunck Henrik and Anind, 2015) contains data recorded from 9 participants performing 6 activities (Biking, Sitting, Standing, Stair down, Stair up and Walking). The participants were given 8 smartphones and 4 smartwatches. For this study, we have used the accelerometer readings from the smartwatch worn on one hand of the user. The accelerometer readings were sampled at a frequency of 50-200 Hz depending on the device used. We have considered “Standing” and “Stair up” as our activity classes and subjects 3 and 9 as our test subjects while considering all other remaining participants across all activities for the training set.
@@ 5. Phase 2: Encoding-based In-context Learning
Figure 6. The broad overview of the setup that the designed study investigates.
@@ 5.3. Annotation as Retrieval of Labels using Examples
Figure 8. t-SNE plots of TFC embeddings for the HHAR (Blunck Henrik and Anind, 2015) across – (a) time domain, (b) frequency domain, and (c) both time and frequency domain embeddings concatenated.
@@ 6.2. Redesigning the Prompt
The following given embeddings correspond to “class 1”: $[[\mathbb{E}^{1}_{11},\ldots,\mathbb{E}^{1}_{u1}]\ldots[\mathbb{E}^{1}_{1v},\ldots,\mathbb{E}^{1}_{uv}]]$ and the following given embeddings correspond to “class $[[\mathbb{E}^{2}_{11},\ldots,\mathbb{E}^{2}_{u1}]\ldots[\mathbb{E}^{2}_{1v},\ldots,\mathbb{E}^{2}_{uv}]]$ classify the embedding $[\mathbb{E}^{q}_{11},\ldots,\mathbb{E}^{q}_{1u}]$ as either “class 1” or “class 2” considering the minimum distance to the example embeddings provided that the distance metric chosen is euclidean distance.
@@ 6.4.3. Impact of Number of Examples
Response
The embedding [ 14.239516 , 7.1703763 ] is classified as “walking_upstairs” based on the minimum distance to the example embeddings, using the Euclidean distance metric.
@@ 7. Discussion and Future Work
    Beyond the Binary Setup: Although in this paper we restricted the number of classes to be annotated by the LLM to two, this setup is completely generalized with the SSL-based pre-trained encoder having no explicit dependency on the number of classes. However, we envision that one of the critical restrictions, with the given assessment using the paid API access of GPT-4, would be in the overall cost of annotating the samples. For example, with the number of classes increasing beyond two, the number of examples provided in the prompt will also increase. This, in general, will increase the number of tokens, which in turn will increase the cost significantly.

### v2

@@ Abstract.
Traditional human-in-the-loop-based annotation for time-series data like inertial data often requires access to alternate modalities like video or audio from the environment. These alternate sources provide the necessary information to the human annotator, as the raw numeric data is often too obfuscated even for an expert. However, this traditional approach has many concerns surrounding overall cost, efficiency, storage of additional modalities, time, scalability, and privacy. Interestingly, recent large language models (LLMs) are also trained with vast amounts of publicly available alphanumeric data, which allows them to comprehend and perform well on tasks beyond natural language processing. Naturally, this opens up a potential avenue to explore the opportunities in using these LLMs as virtual annotators where the LLMs will be directly provided the raw sensor data for annotation instead of relying on any alternate modality. Naturally, this could mitigate the problems of the traditional human-in-the-loop approach. Motivated by this observation, we perform a detailed study in this paper to assess whether the state-of-the-art (SOTA) LLMs can be used as virtual annotators for labeling time-series physical sensing data. To perform this in a principled manner, we segregate the study into two major phases. In the first phase, we investigate the challenges an LLM like GPT-4 faces in comprehending raw sensor data. Considering the observations from phase 1, in the next phase, we investigate the possibility of encoding the raw sensor data using SOTA SSL approaches and utilizing the projected time-series data to get annotations from the LLM. Detailed evaluation with four benchmark HAR datasets shows that SSL-based encoding and metric-based guidance allow the LLM to make more reasonable decisions and provide accurate annotations without requiring computationally expensive fine-tuning or sophisticated prompt engineering.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXX
@@ 1. Introduction
Interestingly, most recent large language models (LLMs) have also been trained with vast knowledge from publicly available data, potentially containing information from numeric data beyond the usual text data (Liu et al., 2023; Yao et al., 2022). This, in turn, opens up an avenue to investigate whether these LLMs can directly annotate the raw sensor data without using alternate modalities. To assess this, we first set up a detailed pilot study (Section 4) whereby we analyze whether the state-of-the-art (SOTA) LLMs like GPT-4 can quickly identify the labels (the activity classes) in a multi-class classification setup. This study shows that although GPT-4 can understand the accelerometer signatures, it fails to separate them across different classes even with a fixed number of bootstrap samples as example observations. Notably, some recent studies like (Liu et al., 2023) have already attempted to investigate how LLMs can be used with sensor data streams and reached similar conclusions.
@@
    Extending LLM’s Capability as Virtual Annotators of the Physical Sensing Data: Based on the findings from the pilot study, we develop an approach to encode the locomotive data in an embedding space that inherently learns the signatures of different classes and thus provides rich contextual information to the LLMs for annotating the accelerometer signatures with the latent embedding of the raw locomotive data. We explore self-supervised contrastive learning with two different setups for this purpose – (a) with time-domain signatures only and (b) with both time and frequency-domain signatures. To the best of our knowledge, this is the first work that helps LLMs enhance their capability for annotating locomotive data from the raw accelerometer with the corresponding human mobility classes without requiring any retraining.
@@
    Thorough Evaluation, Analysis, and Discussion: We perform a detailed investigation using four benchmark HAR datasets on how the encoded accelerometer data offers better reasoning for the LLM to provide accurate annotations with the increasing number of examples. Notably, with paid LLMs like ChatGPT-4, a mere performance assessment with accuracy as the key highlight is not sufficient as there are factors like the number of tokens and rate limits are restrictions that impact the final usage of such a service for annotating a dataset. Understanding this, we in this paper, also assess the time and cost of analysis, summarise the trade-offs, and discuss the advantages of using LLMs as a virtual annotator for time-series physical sensing data. Additionally, we also investigate the inherent limitations of using an unimodal sensor data from a triaxial accelerometer with a comprehensive analysis considering more than two classes.
@@ 2.1. Using LLMs for Understanding Time-Series Data
Interestingly, works like (Liu et al., 2023) used the raw sensor data as an example; however, the idea did not show consistent performance with increasing classes. More recent works, like (Brown et al., 2020; Kim et al., 2024; Xue et al., 2022; Xu et al., 2023a) have attempted to provide more guidance and examples through in-context learning which allows the inclusion of more information within the prompt and does not necessarily need any fine-tuning of the LLM parameters. This approach also has the immense benefit of allowing the transfer of human expertise to the LLM (Xu et al., 2023a; Xie and Min, 2022). Furthermore, works like (Kim et al., 2024; Xu et al., 2023a) also show the importance of providing signal-processing (Xu et al., 2023a) or interpretation-based guidance (Kim et al., 2024) which allows the model to make predictions regarding the query sample. However, one key aspect that we observe in most of these aforementioned works is that in most cases, the input time-series is from specialized sensors like medical (Kim et al., 2024) or temperature sensors (Xu et al., 2023a) for which expert guidance is possible. For example, the prompt can provide what a typical high heart rate means and how it can be observed in the ECG signal. However, with more generalized sensors like accelerometers, it is hardly possible to provide some expert guidance directly, given the inherent nature of the sensor data itself. To mitigate this, works like (Jin et al., 2024) have discussed the idea of using time-series encoders, albeit with a proposal to align the time-series features with language-specific representations, which in turn would need further fine-tuning of the LLM. In this study, however, we explore using a time-series encoder to provide more context to the LLM without fine-tuning. Based on this objective, we assess whether an LLM like GPT-4 can be used as a service to obtain annotations for the unlabeled accelerometer data for HAR.
@@ 3. Datasets Used for the Analysis
UCI HAR (Reyes-Ortiz Jorge, Anguita Davide, Ghio Alessandro, Oneto Luca, Parra Xavier, 2012) 164718 47033 Waist Standing & Walk upstairs 50 Hz
HHAR (Blunck Henrik, Bhattacharya Sourav, Prentow Thor, Kjrgaard Mikkel, and Dey Anind, 2015) 3020605 521516 Arm Standing & Stairs up 50$\sim$200 Hz
@@
    UCI HAR – The UCI HAR dataset (Reyes-Ortiz Jorge, Anguita Davide, Ghio Alessandro, Oneto Luca, Parra Xavier, 2012) contains inertial data for 6 different activities (Laying, Sitting, Standing, Walking, Walking downstairs and Walking upstairs). These activities were performed by wearing a smartphone around their waist. The IMUs have a sampling rate of 50Hz. For our setup, we have chosen “Standing” and “Walking upstairs” as the activity classes and subjects 1, 3, 5, 7, 9, 11, 13, 15 and 17 as the test subjects whereas, all other remaining subjects across all activity classes are part of the training set.
@@
    HHAR – The HHAR dataset (Blunck Henrik, Bhattacharya Sourav, Prentow Thor, Kjrgaard Mikkel, and Dey Anind, 2015) contains data recorded from 9 participants performing 6 activities (Biking, Sitting, Standing, Stair down, Stair up and Walking). The participants were given 8 smartphones and 4 smartwatches. For this study, we have used the accelerometer readings from the smartwatch worn on one hand of the user. The accelerometer readings were sampled at a frequency of 50-200 Hz depending on the device used. We have considered “Standing” and “Stair up” as our activity classes and subjects 3 and 9 as our test subjects while considering all other remaining participants across all activities for the training set.
@@ 5. Phase 2: Encoding-based In-context Learning
Figure 6. The broad overview of the setup that the designed study investigates. This shows an example scenario with data from two classes presented for the task of annotation.
@@ 5.3. Annotation as Retrieval of Labels using Examples
Figure 8. t-SNE plots of (a) SimCLR and (b,c,d) TFC embeddings for the HHAR dataset (Blunck Henrik, Bhattacharya Sourav, Prentow Thor, Kjrgaard Mikkel, and Dey Anind, 2015). Here, (b) time domain, (c) frequency domain, and (d) both time and frequency domain embeddings concatenated obtained from TFC-based encoder.
@@ 6.2. Redesigning the Prompt
The following given embeddings correspond to “class 1”: $[[\mathbb{E}^{1}_{11},\ldots,\mathbb{E}^{1}_{u1}]\ldots[\mathbb{E}^{1}_{1v},\ldots,\mathbb{E}^{1}_{uv}]]$, the following given embeddings correspond to “class 2”: $[[\mathbb{E}^{2}_{11},\ldots,\mathbb{E}^{2}_{u1}]\ldots[\mathbb{E}^{2}_{1v},\ldots,\mathbb{E}^{2}_{uv}]]$ and the following given embeddings corresponds to “class n”: $[[\mathbb{E}^{n}_{11},\ldots,\mathbb{E}^{n}_{u1}]\ldots[\mathbb{E}^{n}_{1v},\ldots,\mathbb{E}^{n}_{uv}]]$ classify the embedding $[\mathbb{E}^{q}_{11},\ldots,\mathbb{E}^{q}_{1u}]$ as either “class 1” or “class 2” . . . . . . . or “class n” considering the minimum distance to the example embeddings provided that the distance metric chosen is euclidean distance.
@@ 6.4. Performance Accuracy
Figure 12. Accuracy of annotations with increasing number of classes with embeddings generated using – (a) SimCLR and (b) TFC. These results show that the accuracy drops with increasing number of classes irrespective of the contrastive setup used.
(a) Three classes
(b) Four classes
(c) Five classes
(d) Six classes
Figure 13. Confusion matrices for multi-class scenarios with embeddings generated using SimCLR. These results expose the inherent limitation of using triaxial accelerometer data for differentiating closely related classes. For example, we see a high confusion from the annotations provided by the LLM with embeddings from closely related classes like “walking” and “walking upstairs”.
@@ 6.4.3. Impact of Number of Examples
Response The embedding [ 14.239516 , 7.1703763 ] is classified as “walking_upstairs” based on the minimum distance to the example embeddings, using the Euclidean distance metric.
@@ 6.6. Reasoning behind the Responses
Figure 14. t-SNE plots from SimCLR embeddings for the UCI HAR dataset. The embeddings of certain classes, like “walking”, “walking upstairs”, and “walking downstairs”, cluster very closely in the embedding space. This results in a drop in the annotation accuracy provided by GPT-4 for the samples from these closely related classes.
6.7. Comprehensive Multi-class Analysis
After confirming our initial hypothesis with two classes, we next conduct a comprehensive multi-class analysis using the UCI HAR dataset due to its smaller test set size than other datasets. Figure 12 provides a detailed report of the overall accuracy we observe with increasing classes. Moreover, as we increased the number of classes, we clearly observed a decline in performance from the LLM. This highlights the importance of identifying classes prone to confusion for deeper insights into this phenomenon, which we will investigate next. From the confusion matrices (shown in Figure 13), we see that up to three classes, the LLM can still annotate Figure 13(a) with 100% accuracy. However, when we start adding new classes in the task of annotation, we observe a significant decline in the performance (see Figure 13(b)–Figure 13(d)). One of the key reasons behind this is the inherent confusion introduced by the new classes, like “walking”, when similar representations are already present for another closely related class, like “walking upstairs” and “walking downstairs”. Similar observation is also observed for classes like “standing” and “sitting”. This can be linked to the overall representational capability of the raw sensor data from the triaxial accelerometer in differentiating between these classes where we see that embeddings for these classes are closely clustered in the embedding space (Figure 14). Our exploration with both SimCLR and TFC frameworks revealed consistent results across all classes, demonstrating that optimal accuracy was consistently attained with a dimensionality of the test embeddings set to 2.
@@ 6.8. Key Observations
  4. (4)
    Effects of Increasing the Number of Class for HAR Dataset – The Choice of the Classes and the Inherent Limitation of the Modality: The overall representational capability of the data itself limits the approach for annotating the raw sensor data using LLMs. For example, when certain classes are added for the annotation task, the representations obtained using the triaxial accelerometer data become confusing enough for the LLM to clearly differentiate the classes from each other. This highlights that it is not just increasing the class but also the choice of the type of the selected class for annotation and the modality used that impacts the overall accuracy of the annotations obtained from the LLM.
@@ 7. Discussion and Future Work
    Increasing Cost with Multiclass Annotation: One of the critical restrictions, with the given assessment using the paid API access of GPT-4, would be in the overall cost of annotating the samples. For example, with the number of classes increasing beyond two, the number of examples provided in the prompt will also increase. This, in general, will increase the number of tokens, which in turn will increase the cost significantly.